OpenStudy (anonymous):

Let f(x) = x2 + 6 and g(x) = x+8/x. Find ( g o f)(-7)

OpenStudy (anonymous):

A. 384/7 B. 63/55 C. -57/7 D. 295/49

OpenStudy (anonymous):

@tgawade

OpenStudy (anonymous):

(gof)(-7) = g(f(-7)) so first we find f(-7) We have f(x) = x^2 + 6 so just replace x by -7 what do you get f(-7) = -----?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Idk sorry

OpenStudy (anonymous):

Ohh wait

OpenStudy (anonymous):

that is f(-7) = (-7)^2 + 6 = ------?

OpenStudy (anonymous):

-43

OpenStudy (anonymous):

I know sorry I wasn't paying attention

OpenStudy (anonymous):

no please check it , square of any negative number is positive .. So f(-7)= 49 + 6 = ---------?

OpenStudy (anonymous):

55

OpenStudy (anonymous):

Answer is B ?

OpenStudy (anonymous):

Yes so we have f(-7) = 55 since we are finding (gof)(-7) = g(f(-7)) ==> g(55) we know that g(x) = (x + 8)/x So g(55) = (55+8)/55 = ----------?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

:)